OpenStudy (anonymous):

Find the future value and interest earned if $56,780 is invested at 5.3% compounded a. quarterly for 23 quarters b. continuously for 15 yrs

OpenStudy (anonymous):

a. F = 56,780(1 + 0.053/4)^(23) = 76,855.95 b. F = 56,780e^(0.053*15) = 125,735.96

OpenStudy (anonymous):

your totally awesome. thank you